Find the inverse Laplace transform L−1{F(s)} of the given function.

F(s) = 8s^2 − 12s + 144 / s(s^2+36)


Your answer should be a function of t. Enclose arguments of functions in parentheses. For example, sin(2x).

L−1{F(s)}= ___________-
1
Expert's answer
2020-07-29T13:31:00-0400

Using partial fractions for "F(s) = \dfrac{8s^{2}-12s+144}{s(s^{2}+36)}" ,


"\begin{aligned}
\dfrac{8s^{2}-12s+144}{s(s^{2}+36)} &= \dfrac{A}{s}+\dfrac{Bs+C}{s^{2}+36}\\
8s^{2}-12s+144 & = A(s^{2}+36) + (Bs+C)s\\
8s^{2}-12s+144 & =(A+B)s^{2}+Cs+36A
\end{aligned}"

Comparing the coefficients of "s^{2}, s" and constant terms, we get

"\begin{aligned}
A+B &= 8\\
C &= -12\\
A = 4 &\Rightarrow B = 4
\end{aligned}"


Therefore,

"\begin{aligned}
L^{-1}\{F(s)\} &= L^{-1}\left\{\dfrac{4}{s} + \dfrac{4s-12}{s^{2}+36} \right\}\\
&=L^{-1}\left\{\dfrac{4}{s}\right\}+L^{-1}\left\{\dfrac{4s}{s^{2}+36}\right\}-L^{-1}\left\{\dfrac{12}{s^{2}+36}\right\}\\
&=4L^{-1}\left\{\dfrac{1}{s}\right\}+4L^{-1}\left\{\dfrac{s}{s^{2}+36}\right\}-2L^{-1}\left\{\dfrac{6}{s^{2}+36}\right\}\\
L^{-1}\{F(s)\}&=4+4\cos(6t)-2\sin(6t) 
\end{aligned}"